AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 293     Entry time: Sun Feb 27 02:38:52 2022
Author: David, Pierre-Michell 
Subject: Current status 

We have decreased the temperatur> 95K --> 90K. At 95K the target was stable, but let's try to increase the density a bit.

Edit 03:54 : 90K --> 85K

Merger, Tape server (WS3): 1300 kB/s writing ok

Leakage current (WS4): 3.27 uA

DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed

Online monitoring (WS 5): done

Attachment 1: 2022-02-27_03-43-36aida.png  165 kB  | Hide | Hide all
2022-02-27_03-43-36aida.png
Attachment 2: 2022-02-27_03-42-55xyto.png  198 kB  | Hide | Hide all
2022-02-27_03-42-55xyto.png
Attachment 3: 2022-02-27_03-42-14exra.png  154 kB  | Hide | Hide all
2022-02-27_03-42-14exra.png
Attachment 4: 2022-02-27_03-41-49extot.png  165 kB  | Hide | Hide all
2022-02-27_03-41-49extot.png
ELOG V3.1.4-unknown